Part 1: Gain =8, R1=5k, R2=20k, R5=100, and V1=1V. Find Vout/Vin.

Part 2: Let R4=0 and R3=∞. Find Vout/Vin. (Hint: make sure the button called Enable Biased Voltage

Display is depressed) (Another Hint: when a resistance is zero, short it; when a resistance is

infinity, delete it).

Part 3: Let R4=2k and R3=∞. Find Vout/Vin.

Part 4: Let R4=0 and R3=1000. Find Vout/Vin.

Part 5:Let R4=2k and R3=1000. Find Vout/Vin.

R1 Vout G1 5k R4 1k R5 100 V1 R2 1V iv 20k Vin GAIN 8 R3 2k
Answer #1

Part 1: Gain=8 circuit

VQ@1 OPAMP.RATED 101 IO4 102 R1 VEE I03 5V R2

Part 2:

Vout/Vin = 0/1 = 0

Part 3:

Vout/Vin = 3.03/1 = 3.03

Part 4:

Vout/Vin = 0/1 = 0

Part 5:

Vout/Vin = 3.03/1 = 3.03